(Q). The axes of the polariser and analyser are inclined to each other at 60 degrees.If the amplitude of polarised light emergent through analyser is A. The amplitude of unpolarised light incident on polariser is. (a)A\2 (b)A (c)2A (d)2(2)^1\2A

(Q).  The axes of the polariser and analyser are inclined to each other at 60 degrees.If the amplitude of polarised light emergent through analyser is A. The amplitude of unpolarised light incident on polariser is.
  (a)A\2   (b)A   (c)2A   (d)2(2)^1\2A

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
Theta1 = 60°
I = Io /2 cos² 60°
I = Io/8
Io = 8I
But I \alpha A²
Hence
\alpha sqrt (I)
Hence
Ao = sqrt(8) A
Ao = 2 * sqrt(2) * A
Option D is correct.